OpenStudy (anonymous):

Determine whether the sequence converges or diverges. If it converges, find the limit. A sub n = sin2n/1+sqrt n

OpenStudy (amistre64):

is that a sin(2n), or a sin^2 (n)

OpenStudy (anonymous):

sin(2n)

OpenStudy (amistre64):

hmm, any ideas on how to test it?

OpenStudy (amistre64):

sin(2n) 1-sqrt n ------ ------- 1+sqrt n 1-sqrt n sin(2n) sqrt n sin(2n) ----- - ------------ 1- n 1- n

OpenStudy (anonymous):

couldn't you just divide everything by 1+sqrt n? I did -1 is less than equal to sin 2n is less than equal to 1 since sin function oscillates from -1 to 1. Then I ran the limit as n approaches infinity.

OpenStudy (amistre64):

let u = 1-n, n = 1-u sin(2-2u) sqrt n sin(2-2u) -------- - ------------ u u sin(2) cos(2u) + cos(2) sin(2u) sqrt n (sin(2) cos(2u) + cos(2) sin(2u)) ------------------------- - ------------------------------- u u

OpenStudy (amistre64):

youd want to use a conjugate to get rid of the sqrt n part underneath

OpenStudy (anonymous):

oh ok that makes sense.

OpenStudy (amistre64):

sin(2) cos(2u) + cos(2) sin(2u) ------------------------- u 2sin(2) cos(2u) + 2cos(2) sin(2u) ------------------------- 2u cos(2u) sin(2u) 2sin(2) ------- + 2cos(2) -------- 2u 2u --------------------------------------------- the tricky part may be this one if my idea holds sqrt (1-u) (sin(2) cos(2u) + cos(2) sin(2u)) - ----------------------------------- u

OpenStudy (amistre64):

and i see i used a sin(a+b) identity instead of the sin(a-b) so some prudence might be warranted

OpenStudy (anonymous):

ok I see what you did. thank you so much!